OpenStudy (anonymous):

John has taken out a loan for college. He started paying off the loan with a first payment of $100. Each month he pays, he wants to pay back 1.1 times as the amount he paid the month before. Explain to John how to represent his first 20 payments in sigma notation. Then explain how to find the sum of his first 20 payments, using complete sentences. Explain why this series is convergent or divergent.

OpenStudy (anonymous):

I'm rather confused by this math question and what to do here...

OpenStudy (tkhunny):

100 100*1.1 100*1.1^2 100*1.1^3 ... 100*1.1^19 There they are. You do the sigma notation.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Nevermind, I got it. Thank you!

OpenStudy (tkhunny):

Did you succeed in adding them up?

OpenStudy (anonymous):

for the most part, i did.

OpenStudy (tkhunny):

100 + 100*1.1 + 100*1.1^2 + 100*1.1^3 + ... + 100*1.1^19 = 100(1 + 1.1 + 1.1^2 + 1.1^3 + ... + 1.1^19) = \(100\dfrac{1-1.1^{20}}{1-1.1} = 100\cdot\dfrac{5.7275}{0.1} = 5727.50\) You must be familiar with Geometric Series. Very helpful skills.

OpenStudy (anonymous):

@Hero is this series convergent or divergent?

hero (hero):

If the series is convergent, then it approaches a certain number. If it is divergent, then it doesn't.

OpenStudy (anonymous):

ok @Hero thank you so much